Suppose there are 180 million employed people and 20 million unemployed people. 

a) What is the unemployment rate? (5 marks)

b) Suppose that 5 million unemployed people give up their search for jobs and become 

discouraged workers. What is the new official unemployment rate? 


1
Expert's answer
2021-08-16T08:49:00-0400

a.

Using the given information, unemployment rate would be:

"=[\frac{Total \space unemployed}{(total \space unemployed+ total \space employed)}]×100\\

=[\frac{20 \space million}{(20 \space million+180 \space million)}]×100\\

=\frac{20 \space million}{200 \space million}×100\\

=10\%"


b.

If 5 million unemployed people become discouraged, they will be removed from the labor force due to which new unemployment rate would be:

"=[\frac{Total \space unemployed}{(total \space unemployed+ total \space employed)}]×100\\

=[\frac{15 \space million}{(15 \space million+180 \space million)}]×100\\

=\frac{15\space million}{195\space million}×100\\

=7.69\%"